# VHE gamma-ray intranight variability from BL Lacertae during the extreme flaring state of 2022

### 25 Mar 2026 v2

@@ 5 Broadband SEDs
A last IXPE campaign on BL Lac during November 2023 resulted in an extremely high optical polarisation of %, not accompanied however with a high X-ray polarisation, leading to a 3 upper limit of 7.4%. This was observed during a period in which BL Lac was behaving as an LSP. Therefore, the X-ray emission was located in the high-energy SED component. These results strongly favour SSC multi-zone or energy-stratified models (see Zhang et al., 2024; Agudo et al., 2025), leaving little room for hadronic contribution, only in the framework of a hybrid model as shown by a detailed analysis of the broadband emission by Liodakis et al. (2025). However, as presented by these authors, hybrid models expect a hardening of the VHE spectrum that is not observed here in order to reproduce the multi-wavelength polarisation signatures.
@@ 6 Discussion
X-ray spectral trends were also observed in BL Lac prior to the 2022 high state (e.g. Weaver et al., 2020). Through an extensive SED modelling, Sahakyan & Giommi (2022) attributed the synchrotron component emerging in X-rays during periods of high state to a second emission zone located beyond the BLR, different from the one dominating the emission in the optical-UV band. The emission from this second region can produce a shift of the synchrotron peak towards higher frequencies, reaching values of even 10 Hz, well within the high-synchrotron-peaked (HSP) range. Indeed, BL Lac is well known for transitioning between the LSP regime — mainly during low flux states — to an IBL/HBL behaviour during active phases (see e.g. Ackermann et al., 2011; Nilsson et al., 2018). Starting on approximately 2019, the source has been going through several years of very high activity, culminating with its three brightest flares ever recorded in 2021, 2022 and 2024. Therefore, the X-ray spectral variability trends observed in the dataset analysed in this work can be related to the emergence of synchrotron emission from freshly accelerated particles in a secondary emission region along the jet. The shift of the X-ray emission from the high- to the low-energy peak of the SED can also be the reason behind the lower correlation of the hard X-ray and optical flux compared to the soft X-ray one, as reported in Appendix D.2. Furthermore, in our model the emission from the secondary region (blob) dominates also in the HE and VHE -ray bands. Therefore, the same freshly accelerated particles are also responsible for the intense flaring activity detected with Fermi-LAT and LST-1 via SSC, as well as for the observed brightening and shifting of the -ray spectrum towards higher energies (see Sect. 4).
@@ 6.2 Week-scale evolution of the broadband emission
We successfully model the broadband emission and evolution of the flare as a SSC two-zone model considering as well the EC contribution. The flare begins with emission dominated by the slowly evolving core, extending up to X-ray energies. Plasmoids (blobs) can develop within some considerably magnetised current sheets within the core through magnetic reconnection. These blobs become primarily responsible for the gamma-ray emission. At this early stage of the flare, the EC scattering also contributes to the gamma-ray emission. As magnetic reconnection sets in, particles in the blob are accelerated to higher energies, increasing its contribution also in the X-ray band. Our spectral fitting reveals several major ejections of plasmoids/blobs in the emission region, namely, BB4, BB6, BB8, and BB10. During all these episodes, a freshly produced blob dominates the gamma-ray emission, and can also have strong contributions in other wavelengths depending on the blob parameters. All those episodes are followed by a relatively low flux episode, where the newly formed blob cools and loses its energy. Since each blob is produced independently, their parameters are expected to be considerably different. On the other hand, the core evolves slowly, thus its physical parameters do not change much between different episodes.
@@
This layout is in agreement with the correlation analysis performed in Appendix D.2, where the high degree of correlation between different bands can be attributed to the co-spatial nature of the broadband emission, with the exception of the lower correlation found with the radio and VHE bands. For the former, it is known that the inner regions of the jet are highly opaque to radio emission, that is typically assumed to come from an outer part (Max-Moerbeck et al., 2014). Several studies have in fact found long delays of hundreds of days between the radio and multi-wavelength emission of blazars (see e.g. Liodakis et al., 2018). We cannot test this hypothesis here as we focus on a shorter period. Nevertheless, such behaviour could explain the lower correlation of this band. We note, however, that the radio band as a lower observing cadence than that in optical, that can also be biasing the measured correlation. Regarding the latter band, the lower correlation of the VHE emission can be explained by the shorter cooling timescales of the highest-energy electrons responsible for the TeV emission. However, as for the radio band correlation, the lower observing cadence of the VHE data can introduce systematic effects in the correlation analysis, diluting the correlation of the TeV emission.
@@
With the configuration assumed in Sect. 5, the low-energy photons from dusty torus are the ones dominating the EC contribution. This EC emission has been used in the past to model the broadband emission of BL Lac (Khatoon et al., 2024). In the present flare, owing to the very bright state in which the source was observed during most of the period, we find that the EC emission becomes relevant only at the early stages of the outburst, when the source is faintest (see the model for BB1 in Fig. 8). We have also considered the possible emission arising from the interplay between the core and blob regions, this is, low-energy photons from one region that can be IC-scattered by electrons of the other region. However, this was found to be negligible due to the compactness of the blob as a result of the size constraints derived from the VHE minute-scale variability. The absorption was also evaluated and taken into account, but was found to be negligible up to energies TeV, where the sources is not detected significantly and only upper limits are estimated.
@@
Overall, the parameters are consistent with the typical values derived from a large sample of blazars (Ghisellini et al., 2010). The core region is mainly responsible for the emission from radio to UV wavelengths, and contributes significantly to the hard X-ray emission during those periods in which the X-ray emission is found in the transition from the low- to the high-energy peak of the SED. During the early moments of the flare, when the soft X-rays are also found in the IC peak, it dominates both the soft and hard X-ray emission. Changes in the parameters between adjacent BBs are rather small considering the time interval between consecutive blocks is generally longer than the light crossing time of the core, reflecting its slow evolution. On the other hand, during those periods when BL Lac fully transitions from LSP to ISP and the X-ray emission is entirely produced by synchrotron radiation, the low energy emission of the blob is the one becoming dominant in the X-ray band. We note that the values of the blob are on the high end of normal values. Zech & Lemoine (2021) interpreted these high values in other BL Lac objects as a thermal coupling of protons and electrons in a weakly magnetised shock, where protons transfer energy to electrons in the shock layer.
@@
The main driver of the variability on day-to-week timescales is the evolution of the particle population of each region during the flare. Especially noticeable is the electron distribution behaviour of the blob during the brightest nights (BB4 and BB8), when it clearly reaches higher energies, in consistency with the IC peak shift towards higher energies reported in Sect. 4. Moreover, for these two nights, an injection of freshly accelerated particles was also needed to explain the remarkably high -ray emission. The blob is also characterised by a higher magnetisation and higher Doppler boosting than the core, with G, and G, , respectively. This, combined with the equipartition values of the core, close to equipartition but still 1, and its sometimes very hard spectral index, are consistent with a magnetic reconnection scenario (see e.g. Jormanainen et al., 2023), discussed in detail in Sect. 6.3. We note that the equipartition values of the blob are far from 1, with the electron energy density dominating on the order of magnitude of . This is a limitation that is regularly found in these models (see Tavecchio & Ghisellini, 2016). This might arise from assuming a still over-simplified picture of the relativistic jets while trying to capture all the observational features, including, week, day and minute scale variability. Nevertheless, the models presented here are able to accurately describe the evolution of the flare while still providing a reasonable solution to observational features such as multi-wavelength correlations, week-, day- and minute-scale VHE variability.
@@ 6.3 Origin of the minute-scale VHE variability
The very fast minute-scale variability demands a compact region with extreme particle acceleration (the blob in our fitting model) on top of the blazar zone (the core in our model) that governs the day-to-week-scale variability. Such a physical setup is often attributed to the mini-jet scenario under magnetic reconnection (Giannios et al., 2009; Petropoulou et al., 2016). A particularly high value of the blob Lorentz factor is also necessary to avoid the VHE -ray attenuation in the compact blob. The optical depth for VHE -rays is given by (Ackermann et al., 2016), where is the observed VHE photon energy, is the optical luminosity, which is the target photons for VHE -rays, and is the shortest observed variability timescale. We find that the Doppler factor can ensure the compact blob is optically thin to the observed VHE photons. This range of Doppler factor values is typically high with respect to those often observed in blazars. Hence, the mini-jet scenario is particularly relevant here to solve this problem and avoid the VHE -ray attenuation in the compact blob. Under the mini-jet scenario, the terminal bulk Lorentz factor of the plasmoid can reach approximately the magnetisation factor of the reconnection site in the co-moving frame (Sironi et al., 2016). As shown in Table 3, the magnetisation factor in the core, estimated by , is generally . The bulk Lorentz factor of the blob relative to the core can be approximated by (if they are both close to the line of sight), which is comparable to the magnetisation factor. Although our fitting model and the above estimates are rather simplified, they are consistent with the mini-jet in magnetic reconnection scenario (Nalewajko et al., 2011).
@@
A plasmoid merger in magnetic reconnection can be an alternative scenario (Zhang et al., 2022). Previous particle-in-cell (PIC) simulations have shown that the SSC can exhibit much higher and faster variability than the synchrotron and EC, due to the highly concentrated electrons freshly accelerated at the plasmoid merging site. The magnetisation factor therein can drop below 1 even though the reconnection site has overall 1 magnetisation, consistent with the fitting parameters derived for the blob (Zhang et al., 2022). In principle, such collisions and mergers of plasma structures can happen in magnetised turbulence as well (Zhang et al., 2023). However, the very high intrinsic X-ray polarisation (49%, see Peirson et al., 2023) measured just two days after the last flare disfavours a turbulent environment.
@@
In addition to the fast variability, the very hard particle spectrum in the core () also indicates a magnetic reconnection origin. Although we fit the radio to optical spectrum by the synchrotron emission from the core, usually the large-scale jet can have a considerable contribution to the radio emission. Therefore, the radio to optical emission intrinsic to the core in our model may require even harder particle spectrum. It is generally believed that only magnetic reconnection in a highly magnetised environment can result in particle spectrum harder than 2 (Guo et al., 2016, 2021). Interestingly, since the correlated quasi-periodic flux and polarisation variations detected by Jorstad et al. (2022) suggested kink instabilities in BL Lac, and the very high optical polarisation degree found in Agudo et al. (2025) indicates the relaxation of a magnetic spring in the jet, with the fast variability reported in the current paper, it is possible that the strong activities of BL Lac since roughly 2020 are associated with giant eruption of magnetised plasma from the central engine, likely due to the magnetically-arrested-disc jet scenario.
@@
Under this picture, our two-zone model essentially consists of a segment of the kinked jet (the core) and plasmoids (the blob) produced via magnetic reconnection that is triggered by the kinked jet from time to time. Since they are co-spatial and physically connected, all bands above the synchrotron self absorption limit should have some correlations. However, the production of plasmoids and their mergers in reconnection is rather random as shown in numerical simulations (Sironi et al., 2016; Zhang et al., 2022). Hence the VHE emission that is fully dominated by the blob can appear less correlated to the optical emission that is fully dominated by the core. Moreover, the models presented here propose a simplified picture of the magnetic reconnection scenario. As the flare develops, plasmoids grow and can be ejected from the reconnection layer with various physical conditions as suggested by recent PIC simulations (Sironi et al., 2016; Zhang et al., 2024). While a fully numerical treatment of plasma dynamics in reconnection is far beyond the scope of this paper, we model these plasmoids as one blob with highly variable physical parameters as shown in many previous studies (e.g. Christie et al., 2019). We note that the two-zone model has significant parameter degeneracy, thus our fitting parameters may not be unique.
@@
Additionally, we have calculated the blob synchrotron and SSC cooling timescales in the observer frame for the electrons responsible for the GeV-TeV -ray emission. The synchrotron cooling timescales were estimated as (see e.g. Abe et al., 2025)
@@
where is the electron mass, is the Thomson cross-section, is the magnetic field of the blob and is the Lorentz factor of the electrons. The blob SSC cooling timescales in the observer frame were computed as , with expressed in the blob reference frame as (Moderski et al., 2005)
@@
where is the photon energy in the rest frame of the blob in units of the electron rest mass energy, , being the photon frequency in the blob frame, is defined by Eq. (C5) in Moderski et al. (2005), is the synchrotron photon number density in the blob rest frame, computed as in Abe et al. (2025). The blob SSC cooling timescales were found to be shorter than the synchrotron ones, consistent with the Compton dominance of – shown by the blob SED. In Table 3 we report the blob SSC cooling timescales at . For the two nights with intranight variability, our models predict SSC cooling timescales of the order of few minutes for electrons radiating at GeV-TeV, in agreement with the minute scale variability observed. Therefore, with all the considerations mentioned above, one possible scenario is that the observed flares at VHE rays are produced by freshly accelerated particles via magnetic reconnection with a cooling dominated by SSC cooling.
@@
We note that the scenario proposed here is based on a magnetically-dominated region. However, an alternative picture based on a two-zone model with a weakly magnetised plasma can still offer a good fitting to the SED as long as the blob has a higher Lorentz factor than that of the core, and larger than the expected minimum Lorentz factor obtained from the TeV fast variability, estimated above. The most natural way to explain higher Lorentz factor in a small blob compared to the overall blazar zone is magnetic reconnection. Nevertheless, and owing to the high degeneracy of blazar SED models (as commented in Sect. 5), other combinations of parameters can lead to valid representations of the SED and its evolution, possibly relying on non-magnetically-dominated scenarios.

### 07 Apr 2026 v3

@@ 5 Broadband SEDs
A final IXPE campaign on BL Lac during November 2023 resulted in an extremely high optical polarisation of %, not accompanied by high X-ray polarisation, yielding a 3 upper limit of 7.4%. BL Lac behaved as an LSP during this period, placing the X-ray emission in the high-energy SED component. These results strongly favour SSC multi-zone or energy-stratified models (see Zhang et al., 2024; Agudo et al., 2025), leaving little room for a hadronic contribution, except within the framework of a hybrid model, as shown by the detailed broadband analysis of Liodakis et al. (2025). However, as shown by these authors, hybrid models expect a hardening of the VHE spectrum to reproduce the multi-wavelength polarisation signatures, which is not observed in the present data.
@@ 6 Discussion
X-ray spectral trends also appear in BL Lac prior to the 2022 high state (e.g. Weaver et al., 2020). Through extensive SED modelling, Sahakyan and Giommi (2022) attribute the synchrotron component emerging in X-rays during periods of high state to a second emission zone located beyond the BLR, distinct from the region that dominates the emission in the optical-UV band. Emission from this second region can shift the synchrotron peak towards higher frequencies, reaching 10 Hz, well within the high-synchrotron-peaked (HSP) range. BL Lac is well known for transitioning between the LSP regime, primarily during low flux states, to an IBL-HBL behaviour during active phases (see e.g. Ackermann et al., 2011; Nilsson et al., 2018). Since approximately 2019, the source has undergone several years of very high activity, culminating in its three brightest flares ever recorded in 2021, 2022, and 2024. Therefore, the X-ray spectral variability trends observed in the dataset analysed here can be attributed to the emergence of synchrotron emission from freshly accelerated particles in a secondary emission region along the jet. The shift of the X-ray emission from the high- to low-energy peak of the SED may also explain the lower correlation between the hard X-ray and optical flux compared to the soft X-ray flux, as reported in Appendix D.2. Furthermore, in our model, the emission from the secondary region (blob) also dominates in the HE and VHE -ray bands. Consequently, these freshly accelerated particles are responsible for the intense flaring activity detected with Fermi-LAT and LST-1 via SSC, as well as for the observed brightening and shifting of the -ray spectrum towards higher energies (see Sect. 4).
@@ 6.2 Week-scale evolution of the broadband emission
We successfully modelled the broadband emission and evolution of the flare as an SSC two-zone model, including the EC contribution. The flare begins with emission dominated by the slowly evolving core, extending up to X-ray energies. Plasmoids (blobs) can develop within highly magnetised current sheets in the core via magnetic reconnection. These blobs are primarily responsible for the gamma-ray emission. At this early stage, EC scattering also contributes to the gamma-ray emission. As magnetic reconnection progresses, particles in the blob are accelerated to higher energies, increasing the blob’s contribution in the X-ray band. Our spectral fitting reveals several major ejections of plasmoids and/or blobs in the emission region, specifically BB4, BB6, BB8, and BB10. During each of these episodes, a newly formed blob dominates the gamma-ray emission and may contribute significantly at other wavelengths, depending on the blob parameters. Each episode is followed by a relatively low-flux period during which the newly formed blob cools and loses its energy. Because each blob forms independently, their parameters are expected to vary considerably, whereas the core evolves slowly, so its physical parameters remain largely unchanged between different episodes.
@@
This scenario is consistent with the correlation analysis presented in Appendix D.2, where the high degree of correlation between different bands is attributed to the co-spatial nature of the broadband emission, with the exception of the lower correlation observed in the radio and VHE bands. For the radio band, it is well established that the inner regions of the jet are highly opaque to radio emission, which typically originates from an outer part (Max-Moerbeck et al., 2014). Several studies report long delays of hundreds of days between the radio and multi-wavelength emission of blazars (see e.g. Liodakis et al., 2018). We could not test this hypothesis as we focused on a shorter period. Nevertheless, such behaviour could explain the lower correlation observed in this band. We note, however, that the radio band has a lower observing cadence than the optical band, which can also bias the measured correlation. For the VHE band, the lower correlation can be explained by the shorter cooling timescales of the highest-energy electrons responsible for the TeV emission. However, as with the radio band, the lower observing cadence of the VHE data may introduce systematic effects in the correlation analysis that dilute the correlation in the TeV emission.
@@
With the configuration adopted in Sect. 5, the low-energy photons from dusty torus dominate the EC contribution. This EC emission has previously been used to model the broadband emission of BL Lac (Khatoon et al., 2024). In the present flare, due to the very bright state of the source during most of the period, we find that the EC emission is only relevant during the early stages of the outburst, when the source is faintest (see the model for BB1 in Fig. 8). We also considered possible emission arising from the interaction between the core and blob regions, that is, low-energy photons from one region being IC-scattered by electrons in the other region. However, this contribution was found to be negligible because of the compactness of the blob as a result of the size constraints derived from the VHE minute-scale variability. We also evaluated the absorption and found it to be negligible up to TeV energies, where the source is not detected significantly and we estimate only upper limits.
@@
Overall, the parameters are consistent with the typical values derived from a large sample of blazars (Ghisellini et al., 2010). The core region primarily drives the emission from radio to UV wavelengths and contributes significantly to the hard X-ray emission during periods where the X-ray emission is transitioning from the low- to high-energy peak of the SED. During the early stages of the flare, when the soft X-rays also lie within the IC peak, the core dominates both the soft and hard X-ray emission. Changes in the parameters between adjacent BBs are relatively small, given that the time interval between consecutive blocks generally exceeds the light-crossing time of the core, reflecting its slow evolution. Conversely, during periods when BL Lac fully transitions from LSP to ISP and the X-ray emission is entirely produced by synchrotron radiation, the low-energy emission from the blob becomes dominant in the X-ray band. The values of the blob are at the high end of normal values. Zech and Lemoine (2021) interpret such high values in other BL Lac objects as resulting from thermal coupling between protons and electrons within a weakly magnetised shock, where protons transfer energy to electrons within the shock layer.
@@
The primary driver of variability on day-to-week timescales is the evolution of the particle populations within each region during the flare. The blob’s electron distribution behaviour of the blob is particularly noticeable during the brightest nights (BB4 and BB8), when it clearly reaches higher energies, consistent with the IC peak shift towards higher energies reported in Sect. 4. Furthermore, for these two nights, an injection of freshly accelerated particles was also necessary to account for the remarkably high -ray emission. The blob is also characterised by higher magnetisation and Doppler boosting than the core, with G, , and G, . This, combined with the equipartition values of the core, which are close to equipartition but still 1, and its occasionally very hard spectral index, is consistent with a magnetic reconnection scenario (see, e.g. Jormanainen et al., 2023), as discussed in detail in Sect. 6.3. The equipartition values of the blob are far from 1, with the electron energy density dominating on the order of magnitude of . This limitation is commonly encountered in such models (see Tavecchio and Ghisellini, 2016). This may result from adopting a an overly simplified representation of relativistic jets while attempting to capture all observational features, including, week-, day-, and minute-scale variability. Nevertheless, the models presented here can accurately describe the evolution of the flare while providing a reasonable solution to observational features such as multi-wavelength correlations and week-, day-, and minute-scale VHE variability.
@@ 6.3 Origin of the minute-scale VHE variability
The extremely fast, minute-scale variability requires a compact region with extreme particle acceleration (the blob in our fitting model) in addition to the blazar zone (the core in our model) , which governs the day-to-week-scale variability. Such a physical setup is often attributed to the mini-jet scenario under magnetic reconnection (Giannios et al., 2009; Petropoulou et al., 2016). The blob also requires a particularly high Lorentz factor to prevent VHE -ray attenuation within the compact blob. The optical depth for VHE -rays is given by (Ackermann et al., 2016), where is the observed VHE photon energy, is the optical luminosity providing the target photons for VHE -rays, and is the shortest observed variability timescale. We find that the Doppler factor ensures that the compact blob is optically thin to the observed VHE photons. This range of Doppler factor values is typically higher than those commonly observed in blazars. Hence, the mini-jet scenario is particularly relevant here to solve this problem and prevent VHE -ray attenuation within the compact blob. In the mini-jet scenario, the terminal bulk Lorentz factor of the plasmoid can approximately equal the magnetisation factor of the reconnection site in the co-moving frame (Sironi et al., 2016). Table 3 shows that the magnetisation factor in the core, estimated by , is generally . The bulk Lorentz factor of the blob relative to the core can be approximated by (assuming both are close to the line of sight), which is comparable to the magnetisation factor. Although our fitting model and the above estimates are rather simplified, they remain consistent with the mini-jet within a magnetic reconnection scenario (Nalewajko et al., 2011).
@@
A plasmoid merger in magnetic reconnection presents an alternative scenario (Zhang et al., 2022). Previous particle-in-cell (PIC) simulations show that the SSC can exhibit much higher and faster variability than the synchrotron and EC, owing to the highly concentrated electrons freshly accelerated at the plasmoid merging site. The magnetisation factor at such sites can drop below one, even though the reconnection site has an overall magnetisation greater than one, consistent with the fitting parameters derived for the blob (Zhang et al., 2022). In principle, such collisions and mergers of plasma structures can also occur in magnetised turbulence (Zhang et al., 2023). However, the very high intrinsic X-ray polarisation (49%, see Peirson et al., 2023), measured just two days after the last flare, disfavours a turbulent environment.
@@
In addition to the rapid variability, the very hard particle spectrum in the core () also indicates a magnetic reconnection origin. Although we fitted the radio-to-optical spectrum by the synchrotron emission from the core, the large-scale jet can often contribute significantly to the radio emission. Consequently, the radio-to-optical emission intrinsic to the core in our model may require an even harder particle spectrum. It is generally accepted that only magnetic reconnection in a highly magnetised environment can produce a particle spectrum harder than two (Guo et al., 2016, 2021). Interestingly, since the correlated quasi-periodic flux and polarisation variations detected by Jorstad et al. (2022) suggest kink instabilities in BL Lac, and the very high optical polarisation degree reported by Agudo et al. (2025) indicates the relaxation of a magnetic spring in the jet, combined with the fast variability observed in this work, it is plausible that the intense activity of BL Lac since approximately 2020 is associated with a giant eruption of magnetised plasma from the central engine, likely arising from a magnetically-arrested-disc jet scenario.
@@
Within this framework, our two-zone model essentially consists of a segment of the kinked jet (the core) and plasmoids (the blob) produced by magnetic reconnection, which is periodically triggered by the kinked jet. As they are co-spatial and physically connected, all bands above the synchrotron self-absorption limit should exhibit some degree of correlation. However, the production of plasmoids and their mergers during reconnection is somewhat random, as demonstrated in numerical simulations (Sironi et al., 2016; Zhang et al., 2022). Consequently, the VHE emission, which is fully dominated by the blob, may appear less correlated with the optical emission, which is fully dominated by the core. Moreover, the models presented here propose a simplified picture of the magnetic reconnection scenario. As the flare develops, plasmoids grow and can be ejected from the reconnection layer under various physical conditions, as indicated by recent PIC simulations (Sironi et al., 2016; Zhang et al., 2024). Although a fully numerical treatment of plasma dynamics during reconnection is beyond the scope of this paper, we modelled these plasmoids as a single blob with highly variable physical parameters, as demonstrated in several previous studies (e.g. Christie et al., 2019). We note that the two-zone model exhibits significant parameter degeneracy; consequently, our fitting parameters may not be unique.
@@
Furthermore, we calculated the synchrotron and SSC cooling timescales of the blob in the observer frame for the electrons responsible for the GeV-TeV -ray emission. We estimated the synchrotron cooling timescales as (see e.g. Abe et al., 2025)
@@
where is the electron mass, is the Thomson cross-section, is the magnetic field of the blob, and is the Lorentz factor of the electrons. We calculated the SSC cooling timescales of the blob in the observer frame as , with expressed in the blob reference frame as (Moderski et al., 2005)
@@
Here, denotes the photon energy in the rest frame of the blob in units of the electron rest-mass energy, , with representing the photon frequency in the blob frame. The function is defined by Eq. (C5) in Moderski et al. (2005), is the synchrotron photon number density in the blob rest frame, calculated as in Abe et al. (2025). We find that the SSC cooling timescales of the blob are shorter than the synchrotron timescales, consistent with the Compton dominance of – inferred from the blob SED. Table 3 lists the SSC cooling timescales of the blob at . For the two nights with intranight variability, our models predict SSC cooling timescales of the order of a few minutes for electrons radiating at GeV-TeV energies, in agreement with the minute-scale variability observed. Therefore, considering all of the above, a plausible scenario is that the observed VHE rays flares are produced by freshly accelerated particles via magnetic reconnection with cooling dominated by SSC processes.
@@
We note that the scenario proposed here assumed a magnetically-dominated region. However, an alternative scenario, based on a two-zone model with a weakly magnetised plasma, can still offer a good fitting to the SED, provided that the blob has a higher Lorentz factor than the core, and exceeds the expected minimum Lorentz factor obtained from the TeV fast variability estimated above. The most natural explanation for a higher Lorentz factor in the small blob relative to the overall blazar zone is magnetic reconnection. Nevertheless, owing to the high degeneracy of blazar SED models (as noted in Sect. 5), other combinations of parameters can yield valid representations of the SED and its evolution, possibly relying on non-magnetically dominated scenarios.